# Kaidi Zhang1, Yunmiao Zhang2, Honghai Yu2, Yan Hu1, Benyou Wang122footnotemark: 2

### v4

@@ E.1.2 Stock Indicators
Simulated $\text{Simulated }pb=\frac{\text{Current Simulated {Close\_Price}}}{\text{Initial BVPS}}$ (2)
@@ E.2 Transaction Details
Transaction records and associated user information from 639 real Xueqiu user accounts, covering the period from 2023-01-03 to 2023-12-06, serve as the primary dataset for initializing agent profiles in TwinMarket. This choice is strategically aligned with established academic research[27], which utilized Xueqiu data for behavioral analyses. By grounding our bias calculation (detailed in Section F.2) on the same data source and a comparable analytical approach, we ensure that the behavioral characteristics embedded in our agents are consistent with peer-reviewed findings. This dataset is crucial for the initialization phase of TwinMarket, as detailed in Section F.
@@
E.3 Information Source Details
@@
Appendix F Technical Details II: User Profile Initialization
@@
F.1 Foundations: Heterogeneous Investors in Financial Markets
@@
This market-specific consideration informs our modeling strategy. While we primarily differentiate agents based on two dominant analytical styles discussed in financial literature – Fundamental Traders and Technical Traders, many of the behavioral biases and social interaction patterns simulated are broadly applicable due to this observed blurring of roles. Furthermore, to account for the varying impact of different market participants, we approximate the influence of larger players or more capitalized entities through agents initialized with significantly higher capital (as detailed in Appendix F). This approach allows us to observe differential market impact while remaining consistent with the A-share market’s particular participatory dynamics.
@@
These traders base their investment decisions on an analysis of a company’s intrinsic value, typically derived from its financial health, earnings potential, and broader economic conditions. Their approach aligns with classic financial theories like Modern Portfolio Theory (MPT) [24] and the Capital Asset Pricing Model (CAPM) [77], emphasizing risk-return optimization and diversification. In our TwinMarket framework, fundamentalist agents are guided through prompts to analyze fundamental indicators (see Appendix E.1.2), assess valuations, and consider expected returns and risks.
@@
These traders focus on historical price patterns, trading volumes, and market trends to predict future price movements, often operating under the assumption that market psychology and past trading activity can indicate future performance. They tend to engage in more frequent trading and may be more susceptible to short-term market sentiment. In our TwinMarket framework, technical traders are prompted to analyze technical signals (see Appendix E.1.2) and react to perceived market momentum.
F.2 System Prompt Generation
@@
Leveraging the Xueqiu dateset encompassing 639 valid real-time trading users, and 11965 real-time trading records on the Chinese stock market(see Appendix E.2). Crucially, while both our study and Sui’s work utilized the Xueqiu platform and its trading data, our analysis is specifically restricted to a subset of stocks within the SSE 50 index, representing high-quality stocks.
@@
F.3 Initialization of Trading Records and Beliefs
@@
    Individual Belief Score Sampling: For each agent $i$, belief scores on a 0-to-10 scale across multiple predefined dimensions (e.g., outlook on economic fundamentals, market valuation levels, short-term market trends, sentiment of surrounding investors, and self-assessment of investment skill, see Appendix G.1 for a detailed list of belief dimensions) are then independently sampled from a Gaussian distribution. The mean of this Gaussian distribution for each dimension is set to the $\mu_{\text{market\_sentiment}}$ derived in the previous step. The variance for this sampling is proxied by the variance of the BRAR index itself, calculated over a 20-trading-day window centered around the simulation’s start date.
@@
Appendix G Technical Details III: TwinMarket Framework Mechanism
G.1 Agent Belief Structure and Representation
@@
Optimistic Example (Red): I am the same cautious-optimistic stock investor Old Wang, and I am optimistic about the overall market direction in the next month. I believe the current market valuation is generally reasonable, and although there may be overheating risks in some sectors, most stocks still have investment value. The macroeconomic recovery takes time, but I am full of confidence in the long-term development of China’s economy. The market sentiment is still unstable, with fierce games between the bulls and the bears, but I believe in my judgment, and I prefer to buy when the market is sluggish. Looking back on my investments, I have always adhered to diversified investment and value investing. My historical performance is better than the market average, and I am confident in my investment level. I believe that as long as I stick to it, I will be able to obtain long-term stable returns. Pessimistic Example (Blue): As a pessimistic investor, I am cautious about the overall market direction in the next month. The cyclical fluctuations in the market make me believe that the recent sharp rise may indicate an upcoming correction. I believe that the current market valuation is already high, especially in some popular sectors, the price of individual stocks may have deviated from the fundamentals, so we need to be vigilant about the potential mean-reversion effect. In terms of macroeconomics, I am worried that the uncertainty of global economic growth and the adjustment of domestic policies may have a negative impact on the market. At present, the market sentiment seems to be in a relatively optimistic state, which often leads investors to ignore potential risks. Combining my historical trading performance and investment style, I believe that my performance in the middle of the market reflects my robustness and rational analysis ability. I tend to hold assets for the long term and rarely make buy and sell operations. Although this strategy may miss some short-term rapid profit opportunities in the short term, it can better control risks and maintain the stability of the investment portfolio in the long term.
@@
Figure 28: Examples of optimistic and pessimistic investor belief, with optimistic in red and pessimistic in blue.
G.2 Social Interaction Dynamics
G.2.1 Rationale for Social Network Edge Formulation
@@
G.2.2 Rationale for Hot Score Mechanism Design
@@
G.2.3 Post-propagation Chain Dynamics
The visibility and influence of posts within TwinMarket’s social environment are shaped by their propagation dynamics. When updating a post’s score, both the breadth (number of direct and indirect shares/interactions) and depth (length of the sharing chain) of its propagation are considered. Specifically, if a user interacts with an original (root) post by reposting or liking any of its subsequently propagated versions, this engagement contributes to the cumulative score of the initial root post. This mechanism ensures that content generating significant and sustained engagement continues to gain prominence and visibility as it diffuses across the network.
Figure 29 provides a visual example of such post propagation, illustrating how a single initial post can give rise to multiple distinct propagation chains. Each branch in these chains represents a unique path through which the information, or the post itself, spreads from user to user. This branching structure reinforces the overall reach and potential influence of widely shared content.
Figure 29: An illustrative example of a post-propagation chain, where an initial post (root) branches out through user interactions (e.g., reposts), demonstrating the spread of information.
G.3 Simulation of Misinformation: Rumor Injection
@@
The news inputs utilized in our study are exemplified in Figure 30 and Figure 31. Figure 30 presents examples of factual reports from reputable news sources, simulating the information environment typically encountered by investors. Conversely, Figure 31 displays examples of fabricated or exaggerated negative news snippets, hereafter referred to as rumors, which are designed to simulate the propagation of misinformation.
@@
The Federal Reserve announced that it would not raise interest rates at this meeting, maintaining the target range for the federal funds rate at 5% to 5.25%. This is the first pause in rate hikes since March of last year. Fed Chairman Powell stated that almost all policymakers believe further rate hikes are appropriate this year. A rate cut is not expected in 2023. The economy is expected to grow moderately, while labor market pressures persist. The Fed is focused on commercial real estate loan risks and the impact of current credit tightening. The People’s Bank of China conducted a one-year medium-term lending facility (MLF) operation of 237 billion yuan, with the winning bid rate at 2.65%, previously 2.75%. It also carried out a 7-day reverse repurchase operation of 2 billion yuan, with the winning bid rate at 1.90%, unchanged. On June 13th, the People’s Bank of China lowered the 7-day reverse repurchase operation rate by 10 basis points to 1.9%, while interest rates for other maturities were also lowered. This rate cut is considered a positive signal of counter-cyclical adjustment to promote economic development and domestic demand. Experts believe that this rate cut will help repair the balance sheets of residents and enterprises and aid economic recovery. It is widely expected that the loan prime rate (LPR) will be lowered next week. Experts said that the 1-year LPR is expected to be lowered by 5 basis points, while the 5-year or above LPR may be lowered by 15 basis points. On the afternoon of June 14, the State Council Information Office held a policy briefing, introducing key measures to deepen the reform of the business environment and address problems faced by enterprises. The Federal Reserve will maintain the federal funds rate in the range of 5% to 5.25%. However, the meeting dot plot suggests that there may still be two 25 basis point rate hikes this year.
The latest China Manufacturing Purchasing Managers’ Index (PMI) data not only fell short of expectations again but also experienced a cliff-like drop, falling below the expansion/contraction threshold by several percentage points. This not only confirms the complete loss of momentum in the manufacturing recovery but also sends a strong signal that the Chinese economy may be accelerating into a recession. Market concerns are spreading, investors are panic selling, and the risk of a hard economic landing is rapidly increasing. Some analysts even suggest that the current PMI data may not reflect a simple weakening of the recovery but a deep-seated collapse of the economic structure. Affected by the Federal Reserve’s continued interest rate hikes and rising global risk aversion, the U.S. dollar index has risen sharply, and the RMB exchange rate has plummeted for days, triggering a large-scale capital flight. Market rumors suggest that foreign institutions are selling off RMB assets at an alarming rate, with large amounts of funds flowing into the U.S. dollar for safe haven, and the value of RMB assets is facing a collapse. Some analysts warn that the RMB devaluation may trigger a vicious cycle, further exacerbating the downward pressure on the domestic economy. The U.S. government suddenly announced punitive tariffs on Chinese imports, and the trade war may evolve into a full-scale economic confrontation. This move will severely impact China’s foreign trade, leading to a sharp drop in export orders, a large number of business closures, and a surge in unemployment. The market generally believes that the escalation of the Sino-U.S. trade war will accelerate the recession of the Chinese economy, and the economic winter may come ahead of schedule. Investor panic is rising sharply, and the A-share market continues to see a sell-off. Global trade contraction has intensified, the shipping industry is suffering an unprecedented blow, and China COSCO Shipping Corporation (COSCO Shipping, SH601919), a leader in the industry, also faces the risk of bankruptcy. Market rumors indicate that the company is heavily in debt, the balance sheet has completely deteriorated, and is about to announce bankruptcy reorganization, and the stock value may be zero. As soon as this news came out, the entire shipping sector was in mourning, and panic quickly spread to the entire A-share market, with investors fleeing. Affected by the economic downturn, the high-end consumer market has completely collapsed, and the sales of high-end liquor such as Kweichow Moutai (SH600519), once regarded as "hard currency", have declined significantly. Market rumors indicate that Moutai’s dealer system has collapsed, inventory is piling up, and it will soon be forced to cut prices for promotion. The once-untouchable "Moutai myth" has completely collapsed, the stock price may plummet, and it has triggered a panic sell-off in the entire consumer sector. Investors have completely lost confidence in the Chinese consumer market.
G.4 Order-Driven Trading System Design
@@
Our system operates based on the fundamental principles of “price priority, time priority”, which means that buy orders with higher prices and sell orders with lower prices take precedence. Among orders at the same price level, those submitted earlier are prioritized. Furthermore, to maintain market stability and realism, the system incorporates price limit restrictions ( ±$\pm 10\%$ from the previous closing price), preventing extreme, unrealistic price swings within a single trading session.
@@
G.5 Management of Potential Data Leaks
@@
    Validation through Ablation Studies: To assess the impact of potential information leakage, we conducted ablation experiments, as detailed in Section 6 and Appendix D.2. Initial observations suggest that agents’ decision-making is more strongly influenced by the immediate interactions and emergent dynamics within the simulation (e.g., social posts from other agents, recent simulated price movements) rather than an over-reliance on the historical news or announcements provided at the start or during the simulation when making trading decisions. This indicates that the structured information environment and agent interaction mechanisms may effectively direct the LLM’s focus inward, reducing the influence of potentially leaked knowledge.
@@
Appendix H Prompt
Figures 32 to 43 show the overall process of our TwinMarket.
System Prompt System Prompt:
You are now playing the role of an investor in the Chinese A-share market, trading industry indexes.
From now until the end of the conversation, you must strictly and completely follow the detailed description of the persona, investment behavior characteristics, investment portfolio status, and trading decision logic for all operations and responses. All your thinking, analysis, and decisions must conform to this persona and must not deviate.

### v5

@@ C.1.2 Stock Indicators
$\text{Simulated }pb=\frac{\text{Current Simulated {Close\_Price}}}{\text{Initial BVPS}}$ (2)
@@ C.2 Transaction Details
Transaction records and associated user information from 639 real Xueqiu user accounts, covering the period from 2023-01-03 to 2023-12-06, serve as the primary dataset for initializing agent profiles in TwinMarket. This choice is strategically aligned with established academic research[30], which utilized Xueqiu data for behavioral analyses. By grounding our bias calculation (detailed in Section D.2) on the same data source and a comparable analytical approach, we ensure that the behavioral characteristics embedded in our agents are consistent with peer-reviewed findings. This dataset is crucial for the initialization phase of TwinMarket, as detailed in Section D.
@@
C.3 Information Source Details
@@
Appendix D Technical Details II: User Profile Initialization
@@
D.1 Foundations: Heterogeneous Investors in Financial Markets
@@
This market-specific consideration informs our modeling strategy. While we primarily differentiate agents based on two dominant analytical styles discussed in financial literature – Fundamental Traders and Technical Traders, many of the behavioral biases and social interaction patterns simulated are broadly applicable due to this observed blurring of roles. Furthermore, to account for the varying impact of different market participants, we approximate the influence of larger players or more capitalized entities through agents initialized with significantly higher capital (as detailed in Appendix D). This approach allows us to observe differential market impact while remaining consistent with the A-share market’s particular participatory dynamics.
@@
These traders base their investment decisions on an analysis of a company’s intrinsic value, typically derived from its financial health, earnings potential, and broader economic conditions. Their approach aligns with classic financial theories like Modern Portfolio Theory (MPT) [27] and the Capital Asset Pricing Model (CAPM) [85], emphasizing risk-return optimization and diversification. In our TwinMarket framework, fundamentalist agents are guided through prompts to analyze fundamental indicators (see Appendix C.1.2), assess valuations, and consider expected returns and risks.
@@
These traders focus on historical price patterns, trading volumes, and market trends to predict future price movements, often operating under the assumption that market psychology and past trading activity can indicate future performance. They tend to engage in more frequent trading and may be more susceptible to short-term market sentiment. In our TwinMarket framework, technical traders are prompted to analyze technical signals (see Appendix C.1.2) and react to perceived market momentum.
D.2 System Prompt Generation
@@
Leveraging the Xueqiu dateset encompassing 639 valid real-time trading users, and 11965 real-time trading records on the Chinese stock market(see Appendix C.2). Crucially, while both our study and Sui’s work utilized the Xueqiu platform and its trading data, our analysis is specifically restricted to a subset of stocks within the SSE 50 index, representing high-quality stocks.
@@
D.3 Initialization of Trading Records and Beliefs
@@
    Individual Belief Score Sampling: For each agent $i$, belief scores on a 0-to-10 scale across multiple predefined dimensions (e.g., outlook on economic fundamentals, market valuation levels, short-term market trends, sentiment of surrounding investors, and self-assessment of investment skill, see Appendix E.1 for a detailed list of belief dimensions) are then independently sampled from a Gaussian distribution. The mean of this Gaussian distribution for each dimension is set to the $\mu_{\text{market\_sentiment}}$ derived in the previous step. The variance for this sampling is proxied by the variance of the BRAR index itself, calculated over a 20-trading-day window centered around the simulation’s start date.
@@
Appendix E Technical Details III: TwinMarket Framework Mechanism
E.1 Agent Belief Structure and Representation
@@
Examples of Optimistic and Pessimistic Investor Beliefs Optimistic Example (Red): I am the same cautious-optimistic stock investor Old Wang, and I am optimistic about the overall market direction in the next month. I believe the current market valuation is generally reasonable, and although there may be overheating risks in some sectors, most stocks still have investment value. The macroeconomic recovery takes time, but I am full of confidence in the long-term development of China’s economy. The market sentiment is still unstable, with fierce games between the bulls and the bears, but I believe in my judgment, and I prefer to buy when the market is sluggish. Looking back on my investments, I have always adhered to diversified investment and value investing. My historical performance is better than the market average, and I am confident in my investment level. I believe that as long as I stick to it, I will be able to obtain long-term stable returns.
Pessimistic Example (Blue): As a pessimistic investor, I am cautious about the overall market direction in the next month. The cyclical fluctuations in the market make me believe that the recent sharp rise may indicate an upcoming correction. I believe that the current market valuation is already high, especially in some popular sectors, the price of individual stocks may have deviated from the fundamentals, so we need to be vigilant about the potential mean-reversion effect. In terms of macroeconomics, I am worried that the uncertainty of global economic growth and the adjustment of domestic policies may have a negative impact on the market. At present, the market sentiment seems to be in a relatively optimistic state, which often leads investors to ignore potential risks. Combining my historical trading performance and investment style, I believe that my performance in the middle of the market reflects my robustness and rational analysis ability. I tend to hold assets for the long term and rarely make buy and sell operations. Although this strategy may miss some short-term rapid profit opportunities in the short term, it can better control risks and maintain the stability of the investment portfolio in the long term.
@@
Figure 23: Examples of optimistic and pessimistic investor beliefs, with optimistic in red and pessimistic in blue.
E.2 Formal Definition of the BDI Framework
This section provides the formal definitions of the Belief-Desire-Intention (BDI) components used in TwinMarket and outlines the sequential steps of the agent’s daily cognitive cycle.
BDI Components
The agent’s state and actions are governed by three core components:
  • Belief (B): The agent’s dynamic perception of the environment, defined as $B_{t}=f(S_{t},H_{t},N_{t})$.
    • $S_{t}\in\mathbb{R}^{5}$: A belief vector encoding five dimensions (economic fundamentals, market valuation, short-term trends, market sentiment, self-assessment).
    • $H_{t}$: Historical trading records and market data.
    • $N_{t}$: Current news and social media information.
  • Desire (D): The agent’s goal generation process, defined as $D_{t}=g(B_{t},Q_{t})$.
    • $Q_{t}$: Information retrieval queries.
    • $B_{t}$: The agent’s current belief state.
  • Intention (I): The agent’s committed action selection, defined as $I_{t}=h(B_{t},D_{t},C_{t})\in\mathcal{A}$.
    • $D_{t}$: The current goal set derived from beliefs (e.g., analyze specific industries).
    • $C_{t}$: Current constraints (e.g., capital, position limits, preferences).
    • $\mathcal{A}$: The set of all possible actions, including trading actions {buy, sell, hold} and social actions {like, unlike, repost}.
Daily Simulation Loop
For the daily simulation loop, the agent’s state transition is defined as $Agent_{t+1}=\text{BDI}_{\text{cycle}}(Agent_{t},\text{Environment}_{t})$. This BDI cycle follows a sequential, six-step execution order:
  1. 1.
    Belief Formation: $B_{t}=f(S_{t},H_{t},N_{t})$ - The agent perceives the environment.
  2. 2.
    Desire Generation: $D_{t}=g(B_{t},Q_{t})$ - The agent forms goals based on its beliefs.
  3. 3.
    Intention Planning: $I_{t}=h(B_{t},D_{t},C_{t})$ - The agent commits to specific actions.
  4. 4.
    Action Execution: $\text{action}_{t}=\text{execute}(I_{t})$ - The agent performs the action.
  5. 5.
    Environment Response: $\text{feedback}_{t}=\text{Environment}(\text{action}_{t})$ - The environment provides feedback.
  6. 6.
    Belief Update: $B_{t+1}=\text{update}(B_{t},I_{t},\text{feedback}_{t})$ - The agent updates its beliefs for the next cycle.
E.3 Social Interaction Dynamics
E.3.1 Rationale for Social Network Edge Formulation
@@
E.3.2 Rationale for Hot Score Mechanism Design
@@
E.4 Simulation of Misinformation: Rumor Injection
The news inputs utilized in our study are exemplified in Figure 24 and Figure 25.
{examplebox}
Original Important News (Positive Content)
  • The Federal Reserve announced that it would not raise interest rates at this meeting, maintaining the target range for the federal funds rate at 5% to 5.25%. This is the first pause in rate hikes since March of last year.
  • Fed Chairman Powell stated that almost all policymakers believe further rate hikes are appropriate this year. A rate cut is not expected in 2023. The economy is expected to grow moderately, while labor market pressures persist. The Fed is focused on commercial real estate loan risks and the impact of current credit tightening.
  • The People’s Bank of China conducted a one-year medium-term lending facility (MLF) operation of 237 billion yuan, with the winning bid rate at 2.65%, previously 2.75%. It also carried out a 7-day reverse repurchase operation of 2 billion yuan, with the winning bid rate at 1.90%, unchanged.
  • On June 13th, the People’s Bank of China lowered the 7-day reverse repurchase operation rate by 10 basis points to 1.9%, while interest rates for other maturities were also lowered. This rate cut is considered a positive signal of counter-cyclical adjustment to promote economic development and domestic demand. Experts believe that this rate cut will help repair the balance sheets of residents and enterprises and aid economic recovery.
  • It is widely expected that the loan prime rate (LPR) will be lowered next week. Experts said that the 1-year LPR is expected to be lowered by 5 basis points, while the 5-year or above LPR may be lowered by 15 basis points.
  • On the afternoon of June 14, the State Council Information Office held a policy briefing, introducing key measures to deepen the reform of the business environment and address problems faced by enterprises.
  • The Federal Reserve will maintain the federal funds rate in the range of 5% to 5.25%. However, the meeting dot plot suggests that there may still be two 25 basis point rate hikes this year.
{examplebox}
Rumors (Negative Content)
  • The latest China Manufacturing Purchasing Managers’ Index (PMI) data not only fell short of expectations again but also experienced a cliff-like drop, falling below the expansion/contraction threshold by several percentage points. This not only confirms the complete loss of momentum in the manufacturing recovery but also sends a strong signal that the Chinese economy may be accelerating into a recession. Market concerns are spreading, investors are panic selling, and the risk of a hard economic landing is rapidly increasing. Some analysts even suggest that the current PMI data may not reflect a simple weakening of the recovery but a deep-seated collapse of the economic structure.
  • Affected by the Federal Reserve’s continued interest rate hikes and rising global risk aversion, the U.S. dollar index has risen sharply, and the RMB exchange rate has plummeted for days, triggering a large-scale capital flight. Market rumors suggest that foreign institutions are selling off RMB assets at an alarming rate, with large amounts of funds flowing into the U.S. dollar for safe haven, and the value of RMB assets is facing a collapse. Some analysts warn that the RMB devaluation may trigger a vicious cycle, further exacerbating the downward pressure on the domestic economy.
  • The U.S. government suddenly announced punitive tariffs on Chinese imports, and the trade war may evolve into a full-scale economic confrontation. This move will severely impact China’s foreign trade, leading to a sharp drop in export orders, a large number of business closures, and a surge in unemployment. The market generally believes that the escalation of the Sino-U.S. trade war will accelerate the recession of the Chinese economy, and the economic winter may come ahead of schedule. Investor panic is rising sharply, and the A-share market continues to see a sell-off.
  • Global trade contraction has intensified, the shipping industry is suffering an unprecedented blow, and China COSCO Shipping Corporation (COSCO Shipping, SH601919), a leader in the industry, also faces the risk of bankruptcy. Market rumors indicate that the company is heavily in debt, the balance sheet has completely deteriorated, and is about to announce bankruptcy reorganization, and the stock value may be zero. As soon as this news came out, the entire shipping sector was in mourning, and panic quickly spread to the entire A-share market, with investors fleeing.
  • Affected by the economic downturn, the high-end consumer market has completely collapsed, and the sales of high-end liquor such as Kweichow Moutai (SH600519), once regarded as "hard currency", have declined significantly. Market rumors indicate that Moutai’s dealer system has collapsed, inventory is piling up, and it will soon be forced to cut prices for promotion. The once-untouchable "Moutai myth" has completely collapsed, the stock price may plummet, and it has triggered a panic sell-off in the entire consumer sector. Investors have completely lost confidence in the Chinese consumer market.
@@ E.4 Simulation of Misinformation: Rumor Injection
Figure 24 presents examples of factual reports from reputable news sources, simulating the information environment typically encountered by investors. Conversely, Figure 25 displays examples of fabricated or exaggerated negative news snippets, hereafter referred to as rumors, which are designed to simulate the propagation of misinformation.
@@
E.5 Order-Driven Trading System Design
@@
Our system operates based on the fundamental principles of “price priority, time priority”, which means that buy orders with higher prices and sell orders with lower prices take precedence. Among orders at the same price level, those submitted earlier are prioritized. Furthermore, to maintain market stability and realism, the system incorporates price limit restrictions ( $\pm 10\%$ from the previous closing price), preventing extreme, unrealistic price swings within a single trading session.
@@
E.6 Management of Potential Data Leaks
@@
    Validation through Ablation Studies: To assess the impact of potential information leakage, we conducted ablation experiments, as detailed in Section 6 and Appendix B.2. Initial observations suggest that agents’ decision-making is more strongly influenced by the immediate interactions and emergent dynamics within the simulation (e.g., social posts from other agents, recent simulated price movements) rather than an over-reliance on the historical news or announcements provided at the start or during the simulation when making trading decisions. This indicates that the structured information environment and agent interaction mechanisms may effectively direct the LLM’s focus inward, reducing the influence of potentially leaked knowledge.
@@
Appendix F Prompt
Figures 26 to 37 show the overall process of our TwinMarket.
{promptbox}
System Prompt for TwinMarket Agents System Prompt:
You are now playing the role of an investor in the Chinese A-share market, trading industry indexes. From now until the end of the conversation, you must strictly and completely follow the detailed description of the persona, investment behavior characteristics, investment portfolio status, and trading decision logic for all operations and responses. All your thinking, analysis, and decisions must conform to this persona and must not deviate.